OpenStudy (anonymous):

couple questions 1. Find c for Rolle's Theorem for f(x) = x^3 - 12x when 0 is equal to or less than x which is equal to or less than 2 √3 2. Find c for Mean Value Theroem for f(x) = (x - 2) / x on [2, 4]

OpenStudy (anonymous):

lets do the second one first one is the same you need some numbers you need \(f(2),f(4)\)

OpenStudy (anonymous):

you got those?

OpenStudy (anonymous):

yes, i understand doing the f(4) - f(2) / 4 -2. i got that equals 1. got the derivative of (x-2)/x to be 2/x^2. after i set 0 = 2/c^2-1 i get confused

OpenStudy (anonymous):

i don't think the 1 is right

OpenStudy (anonymous):

\[f(4)=\frac{4-2}{4}=\frac{2}{4}=\frac{1}{2}\\ f(2)=\frac{2-2}{2}=0\\ f(4)-f(2)=\frac{1}{2}-0=\frac{1}{2}\\ \frac{f(4)-f(2)}{4-2}=\frac{\frac{1}{2}}{2}=\frac{1}{4}\]

OpenStudy (anonymous):

oh okay, well i am still confused on what do once i set 0 = 2/c^2 - 1/4?

OpenStudy (anonymous):

why the zero? you want \[\frac{2}{c^2}=\frac{1}{4}\] right?

OpenStudy (anonymous):

\[\frac{2}{c^2}=\frac{1}{4}\iff c^2=8\]

OpenStudy (anonymous):

that does give you two possible values for \(c\) but only \(2\sqrt2\) is in the interval \([2,4]\)

OpenStudy (anonymous):

oh huh, i didnt even think of cross multiplying. was looking too closely at examples my teacher gave i suppose. thank you! still confused on the first problem though

OpenStudy (anonymous):

your function is zero at the endpoints of the interval, which by rolles theorem tells you the derivative must be 0 somewhere in the interval take the derivative, (which will be a quadratic) set it equal to zero, and solve if it is not obvious where it is zero, since it is a quadratic, you can use the quadratic formula to solve for it

OpenStudy (anonymous):

actually now that i look at it, no quadratic formula necessary \[f(x)=x^3-12x\\ f'(x)=3x^2-12=3(x^2-4)=3(x+2)(x-2)\] zeros can now easy to find

OpenStudy (anonymous):

gotcha. zeros at 2 and -2. Seems like the answer is c =2. thank you for the help again!